OpenStudy (anonymous):

critical point of 4x+4sin(x)

OpenStudy (anonymous):

Derivative set equal to zero.

OpenStudy (anonymous):

4 +4cos(x) =0 cos(x) = -1 x= (2n-1)Pi n= 1,2,......n

OpenStudy (anonymous):

i got zero when i did that but its online homework and its not accepting it

OpenStudy (cwrw238):

differentiating = 4 + 4 cos x at a critical point 4 + 4 cos x = 0 cos x = -1 x is what myko said

OpenStudy (anonymous):

try puting\[(2n-1)\pi\]

OpenStudy (cwrw238):

in degrees x = -180, - 540 and so on

OpenStudy (anonymous):

or plus x = 180, 540

OpenStudy (anonymous):

inflection point would be the second derivative equal to zero

OpenStudy (anonymous):

Answer would be Pi/2

OpenStudy (anonymous):

Cos(x)=-1 X would be 180 degrees

OpenStudy (anonymous):

for the inflection point?

OpenStudy (anonymous):

Second derivative set equal to zero and solve

OpenStudy (anonymous):

is the second derivative -4sin(x)

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Yes, that's correct.

OpenStudy (anonymous):

is the critical point 0?

OpenStudy (anonymous):

Yes that's it